OpenStudy (anonymous):

how to find the perimeter of rhombus given diagonals 30cm ,16cm

OpenStudy (asnaseer):

the sum of te squares of the diagonals in a rhombus is always equal to 4 times the square of one side:\[4s^2=a^2+b^2\]where 's' is the length of one side, 'a' is the length of one diagonal, and 'b' is the length of the other diagonal. once you find 's', the perimeter would just be 4s.

OpenStudy (turingtest):

Here is another way to see itIf the diagonals are a and b we can notice that the diagonals divide the rhombus into 4 congruent right triangles. Each has one leg as a/2 and the other as b/2. We can then use the Pythagorean theorem to find the length of one side\[s^2=(\frac{a}{2})^2+(\frac{b}{2})^2\]Which if you'll notice is the same as asnaseer's formula.